# Heuristic search for linear positive systems

### v1

@@ 3.3 Extension
We have shown above that there exists an SSP formulation in accordance with Definition 3.1 that is equivalent to (1) when $(A+BK)$ is column substochastic for all ${K\in\mathcal{K}}$. By relaxing the first condition of Definition 3.1, allowing $\mathcal{V}^{\prime}$ to contain an infinite number of states, it is possible to construct an equivalent SSP with unit sum transition probabilities, even when column sums of ${(A+BK)}$ exceed unity.
Take an SSP generated by the method described in Section 3.1 that has ${||\mathcal{T}(v,a)||_{1}>1}$ for some $v$, $a$. Let the transition associated with action $a\in\mathcal{A}(v)$ be defined by the probabilities $\{t_{w}(v,a)\}_{w\in\mathcal{V}^{\prime}}$ so that action $a$ gives the probability $t_{w}(v,a)$ of transitioning from state $v$ to state $w$ and $||\mathcal{T}(v,a)||_{1}=\sum_{w\in\mathcal{V}^{\prime}}t_{w}(v,a)$. For each state $v\in\mathcal{V}^{\prime}$ generate a set ${\mathcal{V^{+}}(v)=\{v_{k}\}_{k=1}^{\infty}}$ where ${v_{1}=v}$. Redefine the actions $a\in\mathcal{A}(v)$ for the original states $v\in\mathcal{V}^{\prime}$ so that they fulfill the properties
@@
The condition (5) is equivalent to ${||\mathcal{T}(v,a)||_{1}=1}$. For the states ${v_{k}\in\mathcal{V}^{+}(v)}$, ${k\geq 2}$, define $\mathcal{A}(v_{k})$ by taking each action ${a\in\mathcal{A}(v)}$ and creating a corresponding action ${a^{\prime}\in\mathcal{A}(v_{k})}$ such that
@@
[…]
The infinite size of the state space as alternatively defined in Proposition 3.4 may seem discouraging at first glance. However, the minimization over the action space for the extended states $v_{k}$ corresponding to some $v\in\mathcal{V}^{\prime}$ in the original formulation is identical apart from scaling by a factor for all $k$. This means that the solution to the SSP can still be found by optimization over a finite set of actions, which is guaranteed to contain an optimal solution. Previous results in (Wrobel, 1984) show the possibility of reducing infinite state MDP to a finite "skeleton" under similar conditions. For more discussion on SSP and heuristic search for infinite state spaces, see (Perkins & Barto, 2001) and (Bertsekas, 2018).
Figure 2: Expansion of the state space as described in Section 3.3 to accommodate the unit sum transitions of Definition 3.1, with multiple states $v_{k}$ corresponding to the original $v$. A finite SSP cannot represent the unbounded growth in the state that is possible for, e.g., an unstable system in the formulation (1).
@@ 4 Heuristic bounds
Instances of optimal control problems like (1) may have intractably large state and action spaces, prompting the search for methods that do not require calculations using full dynamics and state. In order to give concrete guarantees on performance, heuristic functions are introduced to bound the optimal cost in any given state. Within the area of control, this has taken the form of relaxed Dynamic Programming, as in (Lincoln & Rantzer, 2006). On the other hand, heuristic search methods that are a direct extension of Dijkstra’s algorithm have been developed for deterministic shortest path problems, an early example being the $A^{*}$ algorithm presented in (Hart et al., 1968). These utilize physical intuition and assumptions about the structure and cost of the analyzed network to construct similar bounding heuristics that guide the search for an optimal solution. For the dynamical systems explored here, the prior knowledge of the system dynamics required to establish such bounds can be represented by some initial stabilizing policy. We first define two functions to bound the optimal cost function given by the solution of (2).
@@
In order to establish an upper heuristic bound in the case of an optimal control problem on the form (1) we also introduce an assumption:
Assumption 3.
A stabilizing control law for a problem on the form (1) is given by $u=\hat{K}x$, where $\hat{K}$ is known.
This allows for an upper bound of the cost in any given state, since following this (likely suboptimal) control law is guaranteed to stabilize the system, bringing the immediate cost $g(x(t),u(t))$ to zero asymptotically. This assumption is necessary to guarantee finite cost in the case of an unstable but stabilizable open-loop system. In the case of Value Iteration for the whole system this would not be a concern; the selected policy generating the minimal cost in each step must be stabilizing. However, the goal of the heuristic search method is to operate only on a portion of the state and action spaces and as such, we cannot guarantee that a stable solution is available in each iteration without Assumption 3. As demonstrated in Section 3.3, the SSP analog of an unstable system requires an infinite state space. As such, similar conditions in the SSP literature, where finite SSP are the main application, are typically lighter and only require reachability of the goal state (point 7 of Definition 3.1).
Remark 4.2.
In principle, the following analysis can be performed using an initial policy that does not stabilize the system in all states. However, as will become clear below, the performance of the resulting methods would be degraded as the resulting upper bound heuristic would contain no quantitative information to guide the search in unstable states. As such, we limit ourselves to the case where Assumption 3 holds.
Given Assumption 3, we next present an initialization of the heuristics, based on the stabilizing policy in the case of the upper bound. The following propositions shows the properties of heuristic bounds for this initial choice, and proves that these hold under Value Iteration.
@@
Consider a system on the form (1) that fulfills Theorem 1. Given Assumption 3 and heuristic functions as in Definition 4.1 with
@@
Iteration of the Bellman equation yields an update on the form (2) as a direct consequence of Theorem 1, since the heuristic function is linear and $\overline{h},\underline{h}\geq 0$. To show that the relation holds after iteration of (2) it is sufficient to note that Value Iteration of $p\leq\overline{h}_{k}$ is a monotone operation with $p$ as a fixed point, together implying that $p\leq\overline{h}_{k+1}$ also holds. An identical argument can be made for the lower bound $\underline{h}$. $\blacksquare$ We have shown that heuristic bounds on the form proposed above indeed bound the optimal cost and can be improved through Value Iteration. However, it is not in general true for an arbitrary heuristic function that the bound in each state improves with every iteration of (2). In the context of SSP, this property holds only for consistent heuristics. The condition required for a lower bound heuristic to be consistent is, as has been noted elsewhere in the literature, equivalent to the Bellman inequality
@@
It is worth noting that this is the exact form of the constraints of the linear program ($iii$) in Theorem 1, giving this program the natural interpretation of a search over consistent lower bound heuristics. For the upper bound heuristic $\overline{h}$, consistency instead corresponds to (10) with the inequality reversed. This property holds for the heuristics given in Proposition 4.3 and is used below to prove the following proposition, guaranteeing monotone improvement under Value Iteration.
@@
Value Iteration of the heuristic bound functions $\overline{H}(x)$ and $\underline{H}(x)$ with initial costs given by Proposition 4.3 yields monotonically improving sequences
$\overline{h}_{0}\geq\overline{h}_{1}\geq\cdots\geq\overline{h}_{k}\;\;\;\textnormal{and}\;\;\;\underline{h}_{0}\leq\underline{h}_{1}\leq\cdots\leq\underline{h}_{k}$
@@
Taking the definition of the initial upper bound heuristic in Proposition 4.3 as $\overline{h}_{0}$ we get
$\overline{h}_{0}$ $=(I-(A+B\hat{K}))^{-\top}(s+\hat{K}^{\top}r)$
$=s+\hat{K}^{\top}r+(A+B\hat{K})^{\top}\overline{h}_{0}$
$\geq s+A^{\top}\overline{h}_{0}+\min_{K\in\mathcal{K}}\left[K^{\top}(r+B^{\top}\overline{h}_{0})\right]$
$=\overline{h}_{1}.$
Since the update (2) is an instance of Value Iteration, which is a monotone operation, $\overline{h}_{0}\geq\overline{h}_{1}$ implies that $\overline{h}_{k}\geq\overline{h}_{k+1}$ for all $k$. For the lower bound we have $s>0$. Further, taking $\underline{h}_{0}=0$ and applying (2) yields
$\underline{h}_{1}$ $=s+A^{\top}\underline{h}_{0}+\min_{K\in\mathcal{K}}\left[K^{\top}(r+B^{\top}\underline{h}_{0})\right]$
@@
which is the value of the initial lower bound given in Proposition 4.3. Hence, $\underline{h}_{0}\leq\underline{h}_{1}$ and the same argument as above gives monotone improvement for all $k$. Finally, strict improvement in at least one element follows directly from the fact that any cost function that is fixed under Value Iteration is also a solution to (2). $\blacksquare$ Going forward, we assume that the value of the heuristic bounds are given a priori, as is customary in the literature on similar methods for SSP. This assumption may be taken to represent some prior theoretical knowledge, or physical intuition gained from practical experience of the system. However, it should be noted that, since the goal of methods using these heuristics is to avoid calculations scaling with the (potentially large) size of the system, it is not a trivial proposition to evaluate the given expression for $\overline{h}$, even when some stabilizing controller $\hat{K}$ is known. The analogue to this problem in the setting of SSP is explored in (McMahan et al., 2005), wherein the authors present a framework for finding a consistent heuristic bound (giving monotone improvement) that scales favorably with the size of the problem.
One possible approach to find $\overline{h}$ given $\hat{K}$ is to leverage any sparsity of the system dynamics to allow for localized calculation of the inverse ${(I-(A+B\hat{K}))^{-1}}$. This can in some cases be artificially enforced by introducing some fictitious local control action $B_{f}$ with cost $r_{f}$ sufficiently high to guarantee that the action will not be part of the optimal solution. In this way we utilize knowledge of the system (knowing a suitable $r_{f}$) that would not ordinarily yield consistency if applied directly, to instead create a consistent heuristic on the form given in Proposition 4.3. For an intuitive example, consider a system of gas pipes where the user desires to find the local optimal routing to dissipate the content on a certain section of the system. Introducing fictitious release valves on all junctures leading from the section of interest renders $A+B\hat{K}$ block diagonal, and the inverse in Proposition 4.3 can thus be calculated on a smaller region of interest for a $\hat{K}$ using the fictitious actions. By selecting the costs $r_{f}$ suitably, given some prior intuition about reasonable optimal costs, we can guarantee that the final solution does not use these actions. The price paid for this enforced consistency of $\overline{h}$ is a deteriorated initial heuristic, since it is based on intentionally costly fictitious actions. As shown below, a less accurate initial heuristic in turn leads to a lower rate of convergence for Value Iteration.
@@
$\underline{h}_{k}\geq\left(1-\frac{1-\delta}{(1-\beta^{-1})^{k}}\right)p.$ (11)
for the resulting heuristic $\underline{h}_{k}$.
@@
The main idea is similar to the proof of convergence as stated in (Lincoln & Rantzer, 2006). The problem (1) has a finite value for any system that fulfills Theorem 1, implying that $\overline{h}<\infty$. This guarantees the existence of $\beta$ and $\delta$ according to the given conditions. Applying the operator (2) to the initial heuristic $\underline{h}_{0}$ gives
$\underline{h}_{1}$ $=s+A^{\top}\underline{h}_{0}+\sum_{i=1}^{M}\min\{r_{i}+B_{i}^{\top}\underline{h}_{0},0\}E_{i}$
@@
The same calculation can be repeated to yield successive bounds up to iteration $k$. An identical bound for convergence of $\overline{h}$ follows from the same reasoning. $\blacksquare$ The bound proven above depends on the global accuracy of the heuristics $\overline{h}$ and $\underline{h}$, indicating that local convergence is slow for an initial policy that performs poorly in some (possibly distant) state. In the following section we present a heuristics-based search algorithm that iteratively calculates local solutions without requirements on global optimality. As a consequence, the above statement can be used to show a faster rate of local convergence when accurate heuristics are available locally.
@@ 5 Heuristic search for positive systems
Next, we propose an algorithm that, starting from some initial state $x_{0}$, locally improves the heuristic bounds in Definition 4.1 by successively expanding a search area $S$. In keeping with the spirit of heuristic search algorithms for SSP, the aim of this method is to find a solution that is sufficiently close to optimality while avoiding an exhaustive search of the full state and action spaces.
Figure 3: Illustration of the search space $S$ in Algorithm 1. Optimal cost functions $\overline{g}_{S}$ and $\underline{g}_{S}$ with regard to the control $K_{S}$ are found using $\overline{h}$ and $\underline{h}$ for states outside $S$. The space is expanded in each iteration to include states with high uncertainty $\overline{h}^{j}-\underline{h}^{j}$ that are most affected under the optimal control laws $\overline{K}_{S}$ and $\underline{K}_{S}$ corresponding to $\overline{g}_{S}$ and $\underline{g}_{S}$.
Let the superscript notation $v^{i}$ for some vector $v$ denote its $i$th element. For the matrices defining the dynamics we define the local dynamics $A_{S}\in\mathbb{R}^{n\times n}$ as the matrix such that the columns with index $i\in\mathcal{V}\setminus S$ are replaced by unit vectors $e_{i}$. Further, define two sets of fully actuating controllers on the states in $S$
$\overline{\mathcal{K}}_{S}$ =$=\{K\in\mathcal{K}:K_{i}=\hat{K}_{i}\;\;\forall i\in\mathcal{V}\setminus S\}$
$\underline{\mathcal{K}}_{S}$ =$=\{K\in\mathcal{K}:K_{i}=\mathbf{0}\;\;\forall i\in\mathcal{V}\setminus S\}.$

### v2

@@ 3.3 Extension
We have shown above that there exists an SSP formulation in accordance with Definition 3.1 that is equivalent to (1) when $(A+BK)$ is column substochastic for all $K\in\mathcal{K}$} (Assumption 2). By relaxing the first condition of Definition 3.1, to allow $\mathcal{V}^{\prime}$ to contain an infinite number of states, it is possible to construct an equivalent SSP with unit sum transition probabilities, even when column sums of (A+BK){(A+BK)} exceed unity.
Take an SSP generated by the method described in Section 3.1 that has $\mathcal{T}(v,a)||_{1}>1$} for some $v$, $a$. Let the transition associated with action $a\in\mathcal{A}(v)$ be defined by the probabilities $\{t_{w}(v,a)\}_{w\in\mathcal{V}^{\prime}}$ so that action $a$ gives the probability $t_{w}(v,a)$ of transitioning from state $v$ to state $w$ and $\mathcal{T}(v,a)||_{1}=\sum_{w\in\mathcal{V}^{\prime}}t_{w}(v,a)$. For each state $v\in\mathcal{V}^{\prime}$ generate a set $\mathcal{V^{+}}(v)=\{v_{k}\}_{k=1}^{\infty}$} where $v_{1}=v$}. Redefine the actions $a\in\mathcal{A}(v)$ for the original states $v\in\mathcal{V}^{\prime}$ so that they fulfill the properties
@@
The condition (7) is equivalent to $\mathcal{T}(v,a)||_{1}=1$}. For the states $v_{k}\in\mathcal{V}^{+}(v)$}, $k\geq 2$}, define $\mathcal{A}(v_{k})$ by taking each action $a\in\mathcal{A}(v)$} and creating a corresponding action $a^{\prime}\in\mathcal{A}(v_{k})$} such that
@@
Following the procedure in Section 3.1 results in the transition $\mathcal{T}(v,a_{1})=\left[\;0\;\;0.8\;\;0.4\;\;0\;\right]^{\top}$} for the first action considered in Example 1. By extending the state space as described in Section 3.3, we instead get transitions defined by $t_{v_{1}}(v_{1},a_{1})=0.2$, $t_{v_{2}}(v_{1},a_{1})=0.4$ and $t_{w_{1}}(v_{1},a_{1})=0.4$ where $v_{k}$ and $w_{k}$ are the states corresponding to the second and third states after expansion. The situation is illustrated in Figure 2.
The infinite size of the state space as required in Proposition 3.4 may seem discouraging at first glance, but is necessary to translate the concept of instability to the setting of SSP. However, the minimization over the action space for the extended states $v_{k}$ corresponding to some $v\in\mathcal{V}^{\prime}$ in the original formulation is identical apart from scaling by a factor for all $k$. This means that the solution to the SSP can still be found by optimization over a finite set of actions, which is guaranteed to contain an optimal solution. Previous results in (Wrobel, 1984) show the possibility of reducing infinite state MDP to a finite "skeleton" under similar conditions. For more discussion on SSP and heuristic search for infinite state spaces, see (Perkins and Barto, 2001) and (Bertsekas, 2018).
Figure 2: Expansion of the state space as described in Section 3.3 to accommodate the unit sum transitions of Definition 3.1, with multiple states $v_{k}$ corresponding to the original $v$. A finite SSP cannot represent the unbounded growth in the state that is possible for, e.g., an unstable system in the formulation (1). We must instead resort to an infinite space of artificial states with nearly identical properties. As is detailed in (Wrobel, 1984), there exist tractable methods for simplification of systems with this structure.
@@ 4 Heuristic bounds
Instances of optimal control problems like (1) may have intractably large state and action spaces, prompting the search for methods that do not require calculations using full dynamics and state. In order to give concrete guarantees on performance, heuristic functions are introduced to bound the optimal cost in any given state. Within the area of control, this has taken the form of relaxed Dynamic Programming, as in (Lincoln and Rantzer, 2006). On the other hand, heuristic search methods that are a direct extension of Dijkstra’s algorithm have been developed for deterministic shortest path problems, an early example being the $A^{*}$ algorithm presented in (Hart et al., 1968). These utilize physical intuition and assumptions about the structure and cost of the analyzed network to construct similar bounding heuristics that guide the search for an optimal solution. For the dynamical systems explored here, the prior knowledge of the system dynamics required to establish such bounds can be represented by some initial stabilizing policy, or a policy stabilizing only a subset of the states. We first define two functions to bound the optimal cost function given by the solution of (3).
@@
4.1 Admissibility
We next present an initialization of the heuristics, based on a stabilizing policy in the case of the upper bound. The following propositions shows that admissibility of the heuristic bounds for this initial choice, and proves that these hold under value iteration.
@@ 4.1 Admissibility
Consider a system on the form (1) that fulfills Theorem 1. Given a stabilizing controller $\hat{K}\in\mathcal{K}$} and heuristic functions as in Definition 4.1 with
@@
Iteration of the Bellman equation yields an update on the form (3) as a direct consequence of Theorem 1, since the heuristic function is linear and $\overline{h},\underline{h}\geq 0$. To show that the relation holds after iteration of (3) it is sufficient to note that value iteration of $p\leq\overline{h}{(k)}$ is a monotone operation with $p$ as a fixed point, together implying that $p\leq\overline{h}{(k+1)}$ also holds. An identical argument can be made for the lower bound $\underline{h}$. $\blacksquare$
4.2 Consistency
We have shown that heuristic bounds on the form proposed above indeed bound the optimal cost and can be improved through value iteration. However, it is not in general true for an arbitrary heuristic function that the bound in each state improves with every fixed-point iteration of (3). In the context of SSP, this property holds only for consistent heuristics. The condition required for a lower bound heuristic to be consistent is, as has been noted elsewhere in the literature, equivalent to the Bellman inequality
@@ 4.2 Consistency
It is worth noting that this is the exact form of the constraints of the linear program (iiiiii) in Theorem 1, giving that program the natural interpretation of a search over consistent lower bound heuristics. For the upper bound heuristic $\overline{h}$, consistency instead corresponds to (12) with the inequality reversed. This property holds for the heuristics given below in Proposition 4.2 and is used to prove monotone improvement under value iteration.
We can establish a finite upper heuristic bound in the case of an optimal control problem on the form (1) if a stabilizing policy is known. This allows for an upper bound of the cost in any given state, since following this (likely suboptimal) policy is guaranteed to stabilize the system, bringing the immediate cost $g(x(t),u(t))$ to zero asymptotically. This is necessary to guarantee finite cost for all possible $x_{0}\in\mathbb{R}^{n}_{+}$ in the case of an unstable but stabilizable open-loop system. In the case of value iteration for the whole system this would not be a concern; the selected policy generating the minimal cost in each step must be stabilizing after a single iteration.
@@
Heuristic bound functions $\underline{H}(x)=s$} and $\overline{H}(x)$ corresponding to the cost under any fixed policy $\hat{K}\in\mathcal{K}$} are consistent, yielding monotonically improving value iteration sequences
$\overline{h}{(0)}\geq\overline{h}{(1)}\geq\cdots\geq\overline{h}{(k)}\;\;\;\textnormal{and}\;\;\;\underline{h}{(0)}\leq\underline{h}{(1)}\leq\cdots\leq\underline{h}{(k)}$
@@
Taking the definition of the initial upper bound heuristic in Proposition 4.2 as $\overline{h}{(0)}$ we get
$\overline{h}{(0)}$ $=(I-(A+B\hat{K}))^{-\top}(s+\hat{K}^{\top}r)$
$=s+\hat{K}^{\top}r+(A+B\hat{K})^{\top}\overline{h}{(0)}$
$\geq s+A^{\top}\overline{h}{(0)}+\min_{K\in\mathcal{K}}\left[K^{\top}(r+B^{\top}\overline{h}{(0)})\right]$
$=\overline{h}{(1)}.$
Since the update (3) is an instance of value iteration, which is a monotone operation, $\overline{h}{(0)}\geq\overline{h}{(1)}$ implies that $\overline{h}{(k)}\geq\overline{h}{(k+1)}$ for all $k$. For the lower bound we have $s>0$. Further, taking $\underline{h}{(0)}=0$ and applying (3) yields
$\underline{h}{(1)}$ $=s+A^{\top}\underline{h}{(0)}+\min_{K\in\mathcal{K}}\left[K^{\top}(r+B^{\top}\underline{h}{(0)})\right]$
@@
which is the value of the initial lower bound given in Proposition 4.2. Hence, $\underline{h}{(0)}\leq\underline{h}{(1)}$ and the same argument as above gives monotone improvement for all $k$. Finally, strict improvement in at least one element follows directly from the fact that any cost function that is fixed under value iteration is also a solution to (3). $\blacksquare$
Going forward, we assume that the value of the heuristic bounds are given a priori, as is customary in the literature on similar methods for SSP, but not necessarily finite in all elements. When one or more states are unstable under an evaluated policy, we set the corresponding elements of the upper bound heuristic to $\overline{h}_{i}=\infty$, to represent the lack of available bounds. This assumption may be taken to represent some prior theoretical knowledge, or physical intuition gained from practical experience of the system. However, it should be noted that, since the goal of methods using these heuristics is to avoid calculations scaling with the (potentially large) size of the system, it is not a trivial proposition to evaluate the given expression for $\overline{h}$ according to Proposition 4.2, even if some global stabilizing controller $\hat{K}$ is known. An analogous problem in the setting of SSP is explored in (McMahan et al., 2005), wherein the authors present a framework for finding and improving a globally consistent heuristic bound (giving monotone improvement) that scales favorably with the size of the problem.
4.3 Superconsistency
Previous heuristic search methods for SSP give guarantees on monotone convergence given a consistent initial heuristic. However, this does not suffice for the upper bound heuristic in the full class of problems (1) unless a stabilizing policy is known. We therefore introduce the notion of superconsistency. Let $E_{ii}$ denote the $i$th diagonal element of $E$. An upper bound heuristic $\overline{h}$ is superconsistent if it satisfies the following inequality:
$\overline{h}\geq s+A^{\top}\overline{h}+\begin{bmatrix}\min\{r_{1}+B_{1}^{\top}\overline{h},0\}E_{11}\\ \vdots\\ \min\{r_{n}+B_{n}^{\top}\overline{h},0\}E_{nn}\end{bmatrix}$} (13)
The rightmost term is the vector corresponding to the sum $\sum_{i=1}^{n}\min\{r_{i}+B_{i}^{\top}p,0\}E_{i}$ in (3) with all off-diagonal elements of $E$ set to zero. As a consequence, it coincides with the normal notion of consistency for $E=I$, but is a stronger requirement in the general case. This property will be essential in proving monotone convergence for a general problem on the form (1) when an initial stabilizing policy is not known.
4.4 Rate of convergence
@@
$\underline{h}{(k)}\geq\left(1-\frac{1-\delta}{(1-\beta^{-1})^{k}}\right)p.$ (14)
for the resulting heuristic $\underline{h}{(k)}$.
@@
The problem (1) has a finite value for any system that fulfills Theorem 1, implying that $\overline{h}<\infty$. This guarantees the existence of $\beta$ and $\delta$ according to the given conditions. Applying the operator (3) to the initial heuristic $\underline{h}{(0)}$ gives
$\underline{h}{(1)}$ $=s+A^{\top}\underline{h}{(0)}+\sum_{i=1}^{M}\min\{r_{i}+B_{i}^{\top}\underline{h}{(0)},0\}E_{i}$
@@
The same calculation can be repeated to yield successive bounds up to iteration $k$. An identical bound for convergence of $\overline{h}$ follows from the same reasoning. $\blacksquare$
The bound proven above depends on the global accuracy of the heuristics $\overline{h}$ and $\underline{h}$, indicating that local convergence is slow for an initial policy that performs poorly in some (possibly distant) state. In the following section we present a heuristics-based search algorithm that iteratively calculates local solutions without requirements on global optimality. As a consequence, the above statement can be used to show a faster rate of local convergence when accurate heuristics are available locally.
@@ 5 Heuristic search for positive systems
Next, we propose an algorithm that, starting from some initial state $x_{0}$, locally improves the heuristic bounds in Definition 4.1 by successively expanding a search area $S$. In keeping with the spirit of heuristic search algorithms for SSP, the aim of this method is to find a solution that is sufficiently close to optimality while avoiding an exhaustive search of the full state and action spaces. The heuristic bounds serve both to measure the distance to optimality and to guide the search, giving faster execution the more detailed the initial information in $\overline{h}$ and $\underline{h}$.
Figure 3: Illustration of the search space $S$ in Algorithm 1. Optimal cost functions $\overline{g}_{S}$ and $\underline{g}_{S}$ with regard to the control $K_{S}$ are found using $\overline{h}$ and $\underline{h}$ for states outside $S$. The space is expanded in each iteration to include states with high uncertainty $\overline{h}_{j}-\underline{h}_{j}$ that are most affected under the optimal control laws $\overline{K}_{S}$ and $\underline{K}_{S}$ corresponding to $\overline{g}_{S}$ and $\underline{g}_{S}$.
5.1 Stabilized systems
In the following setup, we assume knowledge of a stabilizing policy $\hat{K}$ for the full system. This assumption is lifted in the subsequent section. Let the subscript notation $v_{i}$ for vectors $v$ denote its $i$th element, with clarifying parentheses added where subscripts collide (i.e. $(x_{0})_{i}$ indicating the $i$th element of the initial state). For the matrices defining the dynamics we define the local dynamics $A_{S}\in\mathbb{R}^{n\times n}$ as the matrix such that the columns with index $i\in\mathcal{V}\setminus S$ are replaced by unit vectors $e_{i}$. Further, define two sets of fully actuating controllers on the states in $S$:
$\overline{\mathcal{K}}_{S}$ $=\{K\in\mathcal{K}:K_{i}=\hat{K}_{i}\;\;\forall i\in\mathcal{V}\setminus S\}$
$\underline{\mathcal{K}}_{S}$ $=\{K\in\mathcal{K}:K_{i}=\mathbf{0}\;\;\forall i\in\mathcal{V}\setminus S\}.$